# Understanding Adam Requires Better Rotation Dependent Assumptions

### v2

@@ A.2 Reflections and Sampling From The Haar Measure
To sample $R_{n}$ uniformly from $SO(n)$ with respect to the Haar measure, we employ the QR decomposition trick [Mezzadri, 2007, Ozols, 2009], which samples from the Haar measure $\mu$ of the orthogonal group $O(n)$. Let us consider the projection $\pi:O(n)\rightarrow SO(n)$, such that $\pi(\mathbf{R})$ is $\mathbf{R}$ when $\mathbf{R}\in SO(n)$, and $\pi(\mathbf{R})$ simply multiplies the first column of $\mathbf{R}$ by 1-1 when $\mathbf{R}\in O(n)\setminus SO(n)$. The push forward of $\mu$ by $\pi$ is the Haar measure on $SO(n)$. Since Adam is reflection equivariant, rotating with $\pi(\mathbf{R})$ and with $\mathbf{R}$ will lead to identical performance for any $\mathbf{R}\in O(n)$. Thus we can omit to apply $\pi$, and simply sample from $\mu$ using the QR decomposition method.
@@ B.1 Rotations Design Choices
Following [Zhao et al., 2024], this is the only rotation that is dynamic rather than static. Specifically, we compute the full-rank SVD decomposition of the gradient for each layer every 250 steps (recommended frequency in [Zhao et al., 2024]).
@@ B.2 Architectures
We trained a GPT-2 model with 124M parameters on the OpenWebText dataset [Gokaslan and Cohen, 2019] using a configuration designed for efficient pretraining. The model architecture includes 12 layers, 12 attention heads, and a 768-dimensional embedding space, with no bias in LayerNorm or Linear layers. We employed the AdamW optimizer with a peak learning rate of $6\times 10^{-4}$, $\beta_{1}=0.9$, $\beta_{2}=0.95$, and a weight decay of $0.1$, applying gradient clipping of 1.0. Training ran for 100,000 iterations (or 30,000 for some smaller ablations), with learning cosine rate decay starting after a 2,000-iteration warm-up, decaying to a minimum of $6\times 10^{-5}$. We used a sequence length of 1024 and micro batch size of 12 with gradient accumulation steps to simulate an effective batch size of 480 sequences. We additionally tried tuning $\beta_{2}$ by using values $0.9$ and $0.99$. We found that the base AdamW showed slightly better performance, but the globally rotated model’s performance decreased with both of these values, meaning further tuning will not close the gap with the base model. All experiments were performed on four A100 80GB GPUs, leveraging mixed precision. Unless otherwise specified, all optimizer hyperparameters were shared across experiments and set to the default values specified in Karpathy [2022].
@@
We trained a Vision Transformer (ViT) model on the ImageNet-1K dataset [Deng et al., 2009] using the SimpleViT architecture [Beyer et al., 2022]. The model consists of 12 layers, 6 attention heads, a hidden dimension of 384, and an MLP dimension of 1536, with a patch size of 16 and input image size of 224. The AdamW optimizer was employed with a learning rate of 0.001, $\beta_{1}=0.9$, $\beta_{2}=0.999$, $\epsilon=10^{-8}$, and a weight decay of 0.1. We used a cosine learning rate schedule with 5 warm-up epochs. The training was conducted for 100 epochs with a batch size of 1024. All experiments were performed with mixed precision.
@@
We trained a ResNet-50 model [He et al., 2015] on the ImageNet-1K dataset [Deng et al., 2009] using the AdamW optimizer. The optimizer was configured with a learning rate of 0.001, $\beta_{1}=0.9$, $\beta_{2}=0.999$, $\epsilon=10^{-8}$, and a weight decay of 0.0001. We employed a cosine learning rate schedule with 5 warm-up epochs. The training ran for 100 epochs with a batch size of 256.
@@ C.4 Update orthogonality
Additionally, Liu and Su [2025] use a metric inspired by the signal processing literature called the SVD entropy to study Muon updates, which is defined as follows.
@@
Figure 34: The SVD Entropy metric described in [Liu and Su, 2025] throughout training, averaged over network depth. We see similar results to the coefficient of variation, noting that higher is better for this metric.
@@ Appendix D Optimization Algorithms with Rotations
Using the notation of Section˜2, we consider SGD with momentumWe consider the heavy ball formulation here [Polyak, 1964], but the same would hold for Nesterov Accelerated Gradient [Nesterov, 1983]. with learning rate $\eta$, momentum parameter $\beta$, and fixed batches $B_{t}$:
@@ Appendix E SVD Rotations and Muon
Recently, Jordan et al. [2024] proposed Muon, an optimization algorithm for the internal linear layers of neural networks. This algorithm departed from various modifications of Adam on favour of using an “orthogonalized" matrix update. We write the Muon algorithm in Algorithm˜6. We note that a simpler version of the algorithm is described in [Jordan et al., 2024], however, their implementation and the description in subsequent work is as described in Algorithm˜6. We additionally make a notational switch to emphasize that Muon acts only on the matrices of internal layers, Jordan et al. [2024] recommends using a different update scheme (e.g., AdamW) on vector-valued parameters such as bias vectors or LayerNorm parameters (along with the embedding and prediction layer in transformers). We write parameters at time $t$ as $\boldsymbol{\Theta}_{t}$ and their gradients as $\mathbf{G}_{t}$.
@@
While this setting is an oversimplification (the momentum parameters are often crucial for performance), it does offer an interesting connection between Adam’s update in a different basis and more recent algorithms like Muon or Shampoo [Gupta et al., 2018].
@@ Appendix F Common Assumptions in First-Order Optimization Theory
Polyak-Lojasiewicz [Polyak, 1963]
Star-(Strong)-Convexity [Guille-Escuret et al., 2020]
Quadratic Growth [Goujaud et al., 2022]
L-Smoothness ($L_{2}$ norm) [Défossez et al., 2022, Zhou et al., 2024]
Gradient Growth Condition [Zhang et al., 2022]
Bounded Expected Gradient Squared Norm [Zou et al., 2019]
$(L_{0},L_{1})$-Smoothness [Li et al., 2023b]
Restricted Secant Inequality [Guille-Escuret et al., 2022]
Error Bound [Luo and Tseng, 1993, Guille-Escuret et al., 2024]
L-smoothness ($L_{\infty}$ norm)[Guo et al., 2022]
Coordinate-wise $(L_{0},L_{1})$-Smoothness [Crawshaw et al., 2022]
Coordinate-wise “Affine” Variance Noise [Li and Lin, 2024]
Bounded Gradient ($L_{\infty}$) [Reddi et al., 2018]

### v3

@@ A.2 Reflections and Sampling From The Haar Measure
To sample $R_{n}$ uniformly from $SO(n)$ with respect to the Haar measure, we employ the QR decomposition trick [9eafeb2573aa4d7a9d3f0f17ec8c9af5, 2009HowTG], which samples from the Haar measure $\mu$ of the orthogonal group $O(n)$. Let us consider the projection $\pi:O(n)\rightarrow SO(n)$, such that $\pi(\mathbf{R})$ is $\mathbf{R}$ when $\mathbf{R}\in SO(n)$, and $\pi(\mathbf{R})$ simply multiplies the first column of $\mathbf{R}$ by 1-1 when $\mathbf{R}\in O(n)\setminus SO(n)$. The push forward of $\mu$ by $\pi$ is the Haar measure on $SO(n)$. Since Adam is reflection equivariant, rotating with $\pi(\mathbf{R})$ and with $\mathbf{R}$ will lead to identical performance for any $\mathbf{R}\in O(n)$. Thus we can omit to apply $\pi$, and simply sample from $\mu$ using the QR decomposition method.
@@ B.1 Rotations Design Choices
Following [zhao2024galore], this is the only rotation that is dynamic rather than static. Specifically, we compute the full-rank SVD decomposition of the gradient for each layer every 250 steps (recommended frequency in [zhao2024galore]).
@@ B.2 Architectures
We trained a GPT-2 model with 124M parameters on the OpenWebText dataset [Gokaslan2019OpenWeb] using a configuration designed for efficient pretraining. The model architecture includes 12 layers, 12 attention heads, and a 768-dimensional embedding space, with no bias in LayerNorm or Linear layers. We employed the AdamW optimizer with a peak learning rate of $6\times 10^{-4}$, $\beta_{1}=0.9$, $\beta_{2}=0.95$, and a weight decay of $0.1$, applying gradient clipping of 1.0. Training ran for 100,000 iterations (or 30,000 for some smaller ablations), with learning cosine rate decay starting after a 2,000-iteration warm-up, decaying to a minimum of $6\times 10^{-5}$. We used a sequence length of 1024 and micro batch size of 12 with gradient accumulation steps to simulate an effective batch size of 480 sequences. We additionally tried tuning $\beta_{2}$ by using values $0.9$ and $0.99$. We found that the base AdamW showed slightly better performance, but the globally rotated model’s performance decreased with both of these values, meaning further tuning will not close the gap with the base model. All experiments were performed on four A100 80GB GPUs, leveraging mixed precision. Unless otherwise specified, all optimizer hyperparameters were shared across experiments and set to the default values specified in Karpathy2022.
@@
We trained a Vision Transformer (ViT) model on the ImageNet-1K dataset [deng2009imagenet] using the SimpleViT architecture [beyer2022betterplainvitbaselines]. The model consists of 12 layers, 6 attention heads, a hidden dimension of 384, and an MLP dimension of 1536, with a patch size of 16 and input image size of 224. The AdamW optimizer was employed with a learning rate of 0.001, $\beta_{1}=0.9$, $\beta_{2}=0.999$, $\epsilon=10^{-8}$, and a weight decay of 0.1. We used a cosine learning rate schedule with 5 warm-up epochs. The training was conducted for 100 epochs with a batch size of 1024. All experiments were performed with mixed precision.
@@
We trained a ResNet-50 model [he2016deep] on the ImageNet-1K dataset [deng2009imagenet] using the AdamW optimizer. The optimizer was configured with a learning rate of 0.001, $\beta_{1}=0.9$, $\beta_{2}=0.999$, $\epsilon=10^{-8}$, and a weight decay of 0.0001. We employed a cosine learning rate schedule with 5 warm-up epochs. The training ran for 100 epochs with a batch size of 256.
@@ C.4 Update orthogonality
Additionally, liu2025muonscalable use a metric inspired by the signal processing literature called the SVD entropy to study Muon updates, which is defined as follows.
@@
Figure 34: The SVD Entropy metric described in [liu2025muonscalable] throughout training, averaged over network depth. We see similar results to the coefficient of variation, noting that higher is better for this metric.
@@ Appendix D Optimization Algorithms with Rotations
Using the notation of Section˜2, we consider SGD with momentumWe consider the heavy ball formulation here [POLYAK19641], but the same would hold for Nesterov Accelerated Gradient [Nesterov1983AMF]. with learning rate $\eta$, momentum parameter $\beta$, and fixed batches $B_{t}$:
@@ Appendix E SVD Rotations and Muon
Recently, jordan2024muon proposed Muon, an optimization algorithm for the internal linear layers of neural networks. This algorithm departed from various modifications of Adam on favour of using an “orthogonalized" matrix update. We write the Muon algorithm in Algorithm˜6. We note that a simpler version of the algorithm is described in [jordan2024muon], however, their implementation and the description in subsequent work is as described in Algorithm˜6. We additionally make a notational switch to emphasize that Muon acts only on the matrices of internal layers, jordan2024muon recommends using a different update scheme (e.g., AdamW) on vector-valued parameters such as bias vectors or LayerNorm parameters (along with the embedding and prediction layer in transformers). We write parameters at time $t$ as $\boldsymbol{\Theta}_{t}$ and their gradients as $\mathbf{G}_{t}$.
@@
While this setting is an oversimplification (the momentum parameters are often crucial for performance), it does offer an interesting connection between Adam’s update in a different basis and more recent algorithms like Muon or Shampoo [47079].
@@ Appendix F Common Assumptions in First-Order Optimization Theory
Polyak-Lojasiewicz [polyak_gradient_1963]
Star-(Strong)-Convexity [guilleescuret2020studyconditionnumbersfirstorder]
Quadratic Growth [goujaud2022optimalfirstordermethodsconvex]
L-Smoothness ($L_{2}$ norm) [defossez2022a, zhou2024on]
Gradient Growth Condition [NEURIPS2022_b6260ae5]
Bounded Expected Gradient Squared Norm [Zou_2019_CVPR]
$(L_{0},L_{1})$-Smoothness [NEURIPS2023_a3cc5012]
Restricted Secant Inequality [NEURIPS2022_9daab3b4]
Error Bound [luo_error_1993, guilleescuret2023wrongturnssimplegeometry]
L-smoothness ($L_{\infty}$ norm)[guo2022novelconvergenceanalysisalgorithms]
Coordinate-wise $(L_{0},L_{1})$-Smoothness [crawshaw2022]
Coordinate-wise “Affine” Variance Noise [li2024ofracsqrtdt14convergenceratermsprop]
Bounded Gradient ($L_{\infty}$) [j.2018on]